OpenStudy (anonymous):

A coffee blend contains Sumatra beans that cost $5/lb and Kona beans that cost $13/lb. If the blend costs $10/lb, how much of each type of coffee is in 50 lb of the blend

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

So lets define the variables: s = pounds of Sumatra beans k = pounds of Kona beans

TheSmartOne (thesmartone):

We know that s+k = 50 lbs And it says the blend costs $10/lb so then 50 x 10 = $500.

OpenStudy (anonymous):

ok

TheSmartOne (thesmartone):

And we also know that s costs $5 per pounds, and k costs $13 per pound. So we have one equation: \(\sf 5s + 13k = 500\)

TheSmartOne (thesmartone):

And the other equation we got is \(\sf s+k=50\)

TheSmartOne (thesmartone):

So @Amberpatterson just solve these two equations: \(\sf 5s + 13k = 500\) \(\sf s+k=50\)

OpenStudy (anonymous):

ok and then would it be 18.75 ibs of s and 31.25ibs of k

TheSmartOne (thesmartone):

Correct :)

OpenStudy (anonymous):

thank you